OpenStudy (just_one_last_goodbye):

How do I rewrite the rational exponent as a radical by extending the properties of integer exponents.?

OpenStudy (just_one_last_goodbye):

2 7/8 ----- 2 1/4

OpenStudy (just_one_last_goodbye):

@CountryGurl15 can you help me with this one real quick

OpenStudy (anonymous):

lol tbh i have no idea lol sowwy

OpenStudy (just_one_last_goodbye):

esta bien ^_^ gracias de toda forma

OpenStudy (anonymous):

lol si, no problema =)

OpenStudy (just_one_last_goodbye):

@EclipsedStar @Data_LG2 can you guys help me please?

OpenStudy (anonymous):

there are exponents?

OpenStudy (just_one_last_goodbye):

its kinda like fractions on top of each other so yes

OpenStudy (anonymous):

oh you mean like \(\sf \large \frac{2^{7/8}}{2^{1/4}}\) well, you should apply one of the exponent laws first: \(\sf \frac{a^m}{a^n}=a^{m-n}\)

OpenStudy (just_one_last_goodbye):

Yes and ok ^_^ proceed

OpenStudy (anonymous):

so what's 7/8-1/4 ? do that first. what will you get?

OpenStudy (just_one_last_goodbye):

sorry Data was trying to work out other probs ^_^ let me see

OpenStudy (just_one_last_goodbye):

5/8

OpenStudy (anonymous):

right :)